How many grams of water are required to add to 200g of 98% sulfuric acid solution for obtaining a solution with a mass fraction of acid is 20%


1
Expert's answer
2021-03-18T03:01:25-0400

20% = Mass fraction of 1/5 0r 0.5

98% of 200 g sulfuric acid = 196 g

Number and moles of sulphuric is 196/98 = 2 moles

0.5 = 2/ 2+x

1+0.5x = 2

0.5x = 1

X= 2

= 2 moles of water

2×18

= 36 grams of water